Let T : R3-> R3be the linear operator
defined by T(x1, x2, x3) = (x1, x3, -2x2- x3).
Let f(x) = - x³+ 2. Find the operator f(T).
1
Expert's answer
2020-03-26T14:39:36-0400

We have "f(T)=-T^3+2I"

"T\begin{pmatrix}
x_1\\
x_2\\
x_3
\end{pmatrix}=\begin{pmatrix}
1&0&0\\
0&0&1\\
0&-2&-1
\end{pmatrix}\begin{pmatrix}
x_1\\
x_2\\
x_3
\end{pmatrix}"

Then "f(T)\begin{pmatrix}
x_1\\
x_2\\
x_3
\end{pmatrix}=(-T^3+2I)\begin{pmatrix}
x_1\\
x_2\\
x_3
\end{pmatrix}="

"=\left(-\begin{pmatrix}
1&0&0\\
0&0&1\\
0&-2&-1
\end{pmatrix}^3+\begin{pmatrix}
2&0&0\\
0&2&0\\
0&0&2
\end{pmatrix}\right)\begin{pmatrix}
x_1\\
x_2\\
x_3
\end{pmatrix}"

"\begin{pmatrix}
1&0&0\\
0&0&1\\
0&-2&-1
\end{pmatrix}^2=\begin{pmatrix}
1&0&0\\
0&0&1\\
0&-2&-1
\end{pmatrix}\begin{pmatrix}
1&0&0\\
0&0&1\\
0&-2&-1
\end{pmatrix}="

"=\begin{pmatrix}
1&0&0\\
0&-2&-1\\
0&2&-1
\end{pmatrix}"


"\begin{pmatrix}
1&0&0\\
0&0&1\\
0&-2&-1
\end{pmatrix}^3=\begin{pmatrix}
1&0&0\\
0&0&1\\
0&-2&-1
\end{pmatrix}^2\begin{pmatrix}
1&0&0\\
0&0&1\\
0&-2&-1
\end{pmatrix}="

"=\begin{pmatrix}
1&0&0\\
0&-2&-1\\
0&2&-1
\end{pmatrix}\begin{pmatrix}
1&0&0\\
0&0&1\\
0&-2&-1
\end{pmatrix}=\begin{pmatrix}
1&0&0\\
0&2&-1\\
0&2&3
\end{pmatrix}"


"-\begin{pmatrix}
1&0&0\\
0&0&1\\
0&-2&-1
\end{pmatrix}^3+\begin{pmatrix}
2&0&0\\
0&2&0\\
0&0&2
\end{pmatrix}=-\begin{pmatrix}
1&0&0\\
0&2&-1\\
0&2&3
\end{pmatrix}+\begin{pmatrix}
2&0&0\\
0&2&0\\
0&0&2
\end{pmatrix}=\begin{pmatrix}
1&0&0\\
0&0&1\\
0&-2&-1
\end{pmatrix}"

We obtain "f(T)\begin{pmatrix}
x_1\\
x_2\\
x_3
\end{pmatrix}=\begin{pmatrix}
1&0&0\\
0&0&1\\
0&-2&-1
\end{pmatrix}\begin{pmatrix}
x_1\\
x_2\\
x_3
\end{pmatrix}=T\begin{pmatrix}
x_1\\
x_2\\
x_3
\end{pmatrix}" , that is "f(T)=T"

Answer: "f(T)=T"